The Ally Food Corp. has issued 18-year, 8% semi-annual coupon, noncallable bonds at their par value of $1,000 one year ago. Today, the market interest rate on these bonds is 4.5%. What is the current price of the bonds? (Hint: the bonds have 17 years to maturity now).
Answer #1

Coupon = 8%* 1000/2 = 40
Par Value = 1000
Maturity = 17
Number of Periods = 2* 17 = 34
YTM/2 = 4.5%/2 = 2.25%

Price of Bond= PV of Coupons + PV of Par Value = 40*(1-(1+2.25%)-34)/2.25% + 1000/(1+2.25%)34 = 1412.77

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well